# A Mixed-Methods Study on the Implications of Unsafe Rust for Interoperation, Encapsulation, and Tooling

### v3

@@ 6.4. RQ3 - Motivations
Table 7. Survey respondents’ motivations for using unsafe code, as measured by the question shown in Figure 0. All but one participant chose to answer this question. Each cell shows the percentage of participants who identified with both of the motivations for that row and column. The cell at the bottom right hand corner shows the percentage who selected all three motivations.
@@
Table 8. How frequently participants measured the impact of their changes when using unsafe code to improve performance. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in the sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
8a When you choose to use unsafe because it performs faster or is more space efficient, how often do you measure the difference? Used unsafe for performance
[Uncaptioned image]46%36%(76)
[…]
@@
However, a few interview participants used unsafe code to create entire components that were purpose-built to achieve performance. In each situation, performance was seen as a functional requirement for the application domain: “I write a serialization framework…and obviously a goal for it is to be extremely performant” (P0). A participant who took this approach found that it significantly increased the amount of unsafe code in their application, but they felt that this was a reasonable compromise to achieve greater performance. Survey respondents who reported using unsafe code to improve performance did not have a strong tendency toward either large-scale or small-scale use. We found that 26% only pursued “small-scale-optimizations in existing applications,” 20% only built new, “large-scale components,” and 40% used unsafe code for performance in either situation.
Participants did not consistently measure the performance impact of unsafe code. In some situations, they relied on their intuition to determine which design decisions would be best: “I’ll totally admit that…I think this is going to be a hot thing, so I’m going to kind of prematurely optimize” (P0). Survey respondents who used unsafe to increase performance were also inconsistent about measuring its impact. As shown in Table 0, 46% of respondents reported measuring performance at least most of the time, while 36% only sometimes measured performance, if at all. In contrast, Höltervennhoff et al. (Höltervennhoff et al., 2023) found that 6 participants would only use unsafe code to improve performance when the change was “noticeable.” We did not make a distinction between profiling and subjective measures of performance in our survey. However, one of our interview participants performed sophisticated profiling on three versions of a component, each of which had varying amounts of unsafe. The most unsafe-heavy version performed the best, but they opted for one with a moderate amount of unsafe code, balancing safety and performance concerns.
@@
Participants also chose to use unsafe code when they felt it would be easier or more ergonomic than using a safe API. This was the least common motivation cited by interview participants, and only 18% of survey respondents identified with this motivation. Höltervennhoff et al. (Höltervennhoff et al., 2023) indicate that some of their participants would use unsafe code if it “saved them effort”, but it was unclear if this motivation was typical. Our participants usually stated or implied a choice between using safe or unsafe design patterns. For example, one participant chose to use only Rust’s unsafe allocator API rather than a mix of unsafe and safe allocation operations. They found that this approach was easier to reason about, since “if there’s too much abstraction, you lose the information that you need to make it actually sound” (P0). Several participants used the unsafe function transmute to perform unrestricted type conversion. Since types have differing invariants, transmute can only be used in unsafe contexts. One participant used transmutation to simplify the implementation of chess engine, which used three enumerations to represent the position of a chess piece. They found that it was easier to use integer conversion, bitwise operations, and transmutation to convert between these enumerations instead of safely matching on their values. Another participant used transmutation when working with a Rust encapsulation of a C font library. This library exposed C heap allocations as Rust s with short lifetimes—even though these objects would remain valid on the heap for the duration of the program. This participant found it was easier to transmute the Rust objects to have the 'static lifetime, which is indefinite, instead of adapting to the restrictions of the encapsulation.
@@
Table 9. How frequently participants who felt that they had “no clear alternative” to using unsafe code were certain it would be impossible to avoid it. The percentage on the left includes “Always” and “Most of the time”, while the one the right includes “Sometimes” and “Never”. The number of respondents in the sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
9a When you feel that you have no clear alternative other than using unsafe, how often are you certain that it would be completely impossible to accomplish this task using a safe design pattern? Had no alternative to unsafe code
[Uncaptioned image]65%17%(123)
[…]
@@
In most situations, participants perceived that they had “no clear alternative” to using unsafe code. The majority of interview participants and 77% of survey respondents identified with this motivation. Höltervennhoff et al. (Höltervennhoff et al., 2023) observed this to a similar extent; 18 out of their 26 participants claimed to use unsafe code by necessity. Our interview participants often reported being constrained by a combination of Rust’s type system and the properties of their applications. For example, one participant found that it was necessary to implement Send and Sync in a single-threaded application to maintain compatibility with an existing API. The perceived that the implementation was sound by construction. However, each of these traits are declared as unsafe, so they can only be implemented in an unsafe context.
Participants who contributed to just-in-time (JIT) compilers and operating systems found that access memory through raw pointers was unavoidable. One participant found that this was necessary to implement an unwinding table: “I have to jump to that address and I really can’t do anything to verify it…” (P0). In other situations, developers felt that there could be a safe alternative, but they were unaware of how to implement it: “there’s probably a better way that I just don’t know about yet or didn’t know about at the time and I haven’t thought about it” (P0). However, as shown in Table 0, 65% of survey respondents were certain at least most of the time, if not always, that it would be impossible to implement an equivalent safe pattern.
@@
What are Rust developers’ motivations for using unsafe code? Participants were most often motivated to use unsafe code because they felt that there was no safe alternative. In other situations, participants used unsafe code because it performed better or was more ergonomic. For performance, developers pursued both large-scale abstractions and small-scale optimizations, but they did not consistently measure the impact of their changes. Each of these motivations can influence a design decision simultaneously. We expect that interventions that consider the intersection of multiple motivations will be more effective. For example, profiling tools could leverage information on the distribution of unsafe code to assist developers in making a compromise between safety and performance.
@@ 6.5. RQ4 - Encapsulation
Table 10. Survey questions relevant to RQ4. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
10a When you use an unsafe API, how often do you look for documentation to ensure that you meet all of its requirements for safety and correctness? Used unsafe APIs
[Uncaptioned image]85%8%(152)
10b When you expose a safe API for unsafe code, how often do you include runtime checks to ensure that its requirements for correctness and safety are met? Exposed a safe API for unsafe code
[Uncaptioned image]58%29%(141)
[…]
[Uncaptioned image]90%3%(68)
10d When you expose an unsafe API to other users, and it is their responsibility to ensure that certain requirements are met, how often do you include runtime checks for these requirements?
[Uncaptioned image]18%71%(68)
[…]
[Uncaptioned image]6%88%(160)
Regularly wrote unsafe code
[Uncaptioned image]7%85%(85)
[…]
[Uncaptioned image]82%10%(160)
Used unsafe APIs
[Uncaptioned image]83%10%(152)
[…]
@@
Participants reasoned about the soundness and correctness of their unsafe code in terms of invariants. When developing operating systems and embedded applications, these invariants were provided by hardware specifications. For example, one participant who developed an embedded application would write a “magic value” (P0) to RAM to indicate which operation had triggered a reboot. The microcontroller that they used did not erase its memory on reboot, so they could rely on the value being initialized. Some participants perceived hardware-level correctness properties as “orthogonal to Rust” (P0), since they were related to the properties of the language. However, other participants who developed embedded applications relied on third-party libraries to encode hardware requirements into the type system, so that their program would fail to compile if a device was misconfigured.
Developers leveraged Rust’s type system to uphold invariants that they perceived as necessary for soundness. In operating systems, these typically began as operation-specific refinement properties on primitive types, but they quickly built up into system-level invariants; “it gets to a high-level really quickly” (P0). Rust’s aliasing restrictions were a natural fit:
@@
Another participant experimented with using ghost permissions to reason about pointer arithmetic. This method was popularized for Rust by Yanovski et al.’s GhostCell type (Yanovski et al., 2021). Other participants relied on ad-hoc reasoning to ensure that their programs were correct and free of undefined behavior. Some appealed to their prior experience: “[in] C++, you do this all the time” (P0). One participant was accustomed to reasoning about aliasing rules in compiler development, so they felt confident that they were adhering to Rust’s restrictions in unsafe contexts. Other participants validated their decisions by auditing their code, but most had some level of implicit trust; “that’s sort of another… I got to trust it type thing” (P0). This aligns with Höltervennhoff et al. (Höltervennhoff et al., 2023)—a majority of their participants reasoned about unsafe code in terms of contracts or invariants, but some reported using it carelessly, and most felt that it was difficult to write correctly due to its context-sensitivity.
@@
Most survey respondents (61%) predicted that it would be at least somewhat easy for another developer at their skill level to understand a random unsafe block or function from their code. Interview participants who isolated their unsafe code had more confidence that its requirements were met, especially if they went beyond what Rust’s type system could verify. However, participants reported that unsafe code was prevalent in JIT compilers and operating systems. Domain-specific, non-local reasoning was necessary to understand any arbitrary unsafe code snippet: “you have to know the innards of the JIT-compiled function…so it’s just not encapsulated” (P0). Though unsafe code may have been isolated, it is unclear if participants used it minimally. As shown in Table 0, more than 80% (6.5) of survey respondents only sometimes refactored their code to remove unsafe features, if at all. This distribution was similar for the subset of respondents who had regularly written unsafe code.
@@
Most interview participants and 88% of survey respondents reported that they would attempt to encapsulate unsafe code beneath safe APIs. This was also common for Höltervennhoff et al. (Höltervennhoff et al., 2023), who found that 22 of their 26 participants mentioned safe interfacing. Our interview participants would create encapsulations where they perceived that the requirements for their unsafe code were satisfied by either Rust’s type system or what they reasoned to be local invariants. However 43% of our survey respondents reported exposing unsafe APIs, while only 6 of Höltervenhoff et al. (Höltervennhoff et al., 2023)’s 26 participants did so.
Interview participants cited the same motivations for exposing unsafe APIs as they did for using unsafe code in any capacity. One participant exposed unsafe versions of safe API endpoints which had the same behavior, but they did not include runtime checks. Another interview participant exposed unsafe APIs so that users could circumvent their safe encapsulations, in case they became difficult to use. Each of these practices aligns with another participant’s perception of a cultural value among Rust developers to make API surfaces as detailed as possible: “Rust has a tendency to… make the API as complicated as it needs to be to fully represent what is actually happening behind the scenes” (P0). Other participants indicated that they would only expose an unsafe API when it would be impossible to encapsulate without placing the burden of correctness on the user. Our survey respondents also identified with these motivations; 76% exposed unsafe APIs because it would be impossible to encapsulate them without preconditions, while 51% exposed unsafe APIs for performance and 13% did so for ergonomics. Höltervennhoff et al. (Höltervennhoff et al., 2023) also found that participants would expose unsafe APIs by necessity or to improve performance, but it was unclear if ergonomics was also a motivation.
Regardless of motivation, 61% of respondents claimed that they would only expose unsafe APIs when they had requirements beyond what Rust’s type system could verify, and they typically documented these requirements. Results for Question 6.5 indicate that 90% of respondents who exposed unsafe APIs would document their safety requirements most of the time, if not always. However, respondents did not typically add any additional safeguards to prevent unsafe APIs from being used incorrectly, since 71% (6.5) only sometimes inserted run-time checks within these APIs, if at all. This was more common when exposing safe APIs; 58% (6.5) inserted checks at least most of the time when exposing a safe API for unsafe code. When using unsafe APIs, 85% (6.5) of participants would usually look for documentation to inform their design decisions.
@@
An overwhelming majority of survey respondents also preferred safe APIs. More than 80% (6.5) of respondents claimed that they would choose a safe API over an unsafe alternative most of the time, if not always—regardless of whether or not they had ever used an unsafe API before.
Table 11. Survey questions relevant to RQ4. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parenthesis on the righthand side.
{tblr}
width=colspec=Q[c,m]X[4,l,m]X[1,c,m]Q[c,m] ID Question Sample Distribution
11a When you expose a safe API for unsafe code, how often are all of its requirements for correctness and safety satisfied by the properties of Rust’s type system? Exposed a safe API for unsafe code
[Uncaptioned image]62%18%(141)
11b How often is the Rust community’s guidance and documentation adequate for you to know how to use unsafe correctly? All
[Uncaptioned image]70%16%(160)
[…]
@@
Höltervennhoff et al. (Höltervennhoff et al., 2023) had similar findings; 16 of their 26 participants felt some degree of uncertainty about whether their unsafe code was correct. Our interview participants were most often uncertain when unsafe was pervasive or when they called foreign functions. When describing their uncertainty about encapsulation, our participants compared the “theory” of safety to the “practice” of how an API would be used. When these concepts didn’t align, errors occurred. One participant observed this mismatch with a Rust crate that encapsulated a foreign library. The library used a memory-mapped file, and the API made it possible to unmap the file while the user retained a reference to it, leading to a use-after-free error. Another participant struggled with discrepancies between the interface and the implementation of Rust’s Layout API, which describes the size and alignment of an allocation. The design of Layout was changed to “tweak the safety requirements” (P0), which regretfully caused previously correct code to exhibit undefined behavior.
@@
A few interview participants encountered situations where the precise definition of undefined behavior was either unclear or overly restrictive. For example, one participant felt that it would be necessary to create a mutable reference to uninitialized memory. This is considered undefined behavior by Rust’s current semantics (Durham, 2022), but the participant was uncertain about how this pattern would be problematic in practice. Another participant wished to be able to take a reference to the first field of a product type and use it to create valid references to adjacent fields using integer offsets. This was considered undefined behavior under Rust’s Stacked Borrows (Jung et al., 2019) aliasing model, but it is accepted under the newer Tree Borrows model (Villani et al., 2023). Additionally, multiple participants were uncertain about how to handle the case when a program panics while a value is being dropped.
Survey respondents were somewhat more certain about their decisions. As shown in Table 0, more than 60% (6.5) of respondents who had exposed a safe API for unsafe code were certain at least most of the time, if not always, that its safety properties were satisfied by Rust’s type system. Additionally, 70% (6.5) of respondents found that the Rust community’s documentation and guidance was adequate for them to know how to use unsafe code correctly at least most of the time. However, only 6% felt that it was always adequate, suggesting that improved documentation would be helpful for the overwhelming majority of respondents. Similar to the questions in Tables 6.3 and  6.4, we added “Unsure” as a response option for the questions shown in Table 0 since they rely on developers’ subjective notions of certainty and adequacy.

### v4

@@ 5.4. RQ3 - Motivations
Table 6. Survey respondents’ motivations for using unsafe code, as measured by the question shown in Figure 0. All but one participant chose to answer this question. Each cell shows the percentage of participants who identified with both of the motivations for that row and column. All three motivations were selected by 7% of participants.
@@
Table 7. How frequently participants measured the impact of their changes when using unsafe code to improve performance. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in the sample is included in parentheses on the righthand side.
ID Question Sample Distribution
7a When you choose to use unsafe because it performs faster or is more space efficient, how often do you measure the difference? Used unsafe for performance 46%36%(76)
Always    Most of the time    About half the time    Sometimes    Never
@@
However, a few interview participants used unsafe code to create entire components that were purpose-built to achieve performance. In each situation, performance was seen as a functional requirement for the application domain: “I write a serialization framework…and obviously a goal for it is to be extremely performant” (P3). A participant who took this approach found that it significantly increased the amount of unsafe code in their application, but they felt that this was a reasonable compromise to achieve greater performance. Survey respondents who reported using unsafe code to improve performance did not have a strong tendency toward either large-scale or small-scale use. We found that 26% only pursued “small-scale-optimizations in existing applications,” 20% only built new, “large-scale components,” and 42% used unsafe code for performance in either situation.
Participants did not consistently measure the performance impact of unsafe code. In some situations, they relied on their intuition to determine which design decisions would be best: “I’ll totally admit that…I think this is going to be a hot thing, so I’m going to kind of prematurely optimize” (P11). Survey respondents who used unsafe code to increase performance were also inconsistent about measuring its impact. As shown in Table 0, 46% of respondents reported measuring performance at least most of the time, while 36% only sometimes measured performance, if at all. In contrast, Höltervennhoff et al. (Höltervennhoff et al., 2023) found that 6 participants would only use unsafe code to improve performance when the change was “noticeable.” We did not make a distinction between profiling and subjective measures of performance in our survey. However, one of our interview participants performed sophisticated profiling on three versions of a component, each of which had varying amounts of unsafe. The most unsafe-heavy version performed the best, but they opted for one with a moderate amount of unsafe code, balancing safety and performance concerns.
@@
Participants also chose to use unsafe code when they felt it would be easier or more ergonomic than using a safe API. This was the least common motivation cited by interview participants, and only 18% of survey respondents identified with this motivation. Höltervennhoff et al. (Höltervennhoff et al., 2023) indicate that some of their participants would use unsafe code if it “saved them effort”, but it was unclear if this motivation was typical. Our participants usually stated or implied a choice between using safe or unsafe design patterns. For example, one participant chose to use only Rust’s unsafe allocator API rather than a mix of unsafe and safe allocation operations. They found that this approach was easier to reason about, since “if there’s too much abstraction, you lose the information that you need to make it actually sound” (P4). Several participants used the unsafe function transmute to perform unrestricted type conversion. Since types have differing invariants, transmute can only be used in unsafe contexts. One participant used transmutation to simplify the implementation of chess engine, which used three enumerations to represent the position of a chess piece. They found that it was easier to use integer conversion, bitwise operations, and transmutation to convert between these enumerations instead of safely matching on their values. Another participant used transmutation when working with a Rust encapsulation of a C font library. This library exposed C heap allocations to Rust as references with short lifetimes—even though these objects would remain valid on the heap for the duration of the program. This participant found it was easier to transmute the Rust objects to have the 'static lifetime, which is indefinite, instead of adapting to the restrictions of the encapsulation.
@@
Table 8. How frequently participants who felt that they had “no clear alternative” to using unsafe code were certain it would be impossible to avoid it. The percentage on the left includes “Always” and “Most of the time”, while the one the right includes “Sometimes” and “Never”. The number of respondents in the sample is included in parentheses on the righthand side.
ID Question Sample Distribution
8a When you feel that you have no clear alternative other than using unsafe, how often are you certain that it would be completely impossible to accomplish this task using a safe design pattern? Had no alternative to unsafe code 65%17%(123)
Always    Most of the time    About half the time    Unsure    Sometimes    Never
@@
In most situations, participants perceived that they had “no clear alternative” to using unsafe code. The majority of interview participants and 77% of survey respondents identified with this motivation. Höltervennhoff et al. (Höltervennhoff et al., 2023) observed this to a similar extent; 18 out of their 26 participants claimed to use unsafe code by necessity. Our interview participants often reported being constrained by a combination of Rust’s type system and the properties of their applications. For example, one participant found that it was necessary to implement Send and Sync in a single-threaded application to maintain compatibility with an existing API. They perceived that the implementation was sound by construction. However, both of these traits are unsafe and only be implemented in an unsafe context.
Participants who contributed to just-in-time (JIT) compilers and operating systems found it difficult to avoid accessing memory through raw pointers. One participant found that this was necessary to implement an unwinding table: “I have to jump to that address and I really can’t do anything to verify it…” (P12). In other situations, developers felt that there could be a safe alternative, but they were unaware of how to implement it: “there’s probably a better way that I just don’t know about yet or didn’t know about at the time and I haven’t thought about it” (P18). However, as shown in Table 0, 65% of survey respondents were certain at least most of the time, if not always, that it would be impossible to implement an equivalent safe pattern.
@@
What are Rust developers’ motivations for using unsafe code?
Participants were most often motivated to use unsafe code because they felt that there was no safe alternative. In other situations, participants used unsafe code because it performed better or was more ergonomic. For performance, developers pursued both large-scale abstractions and small-scale optimizations, but they did not consistently measure the impact of their changes. Each of these motivations can influence a design decision simultaneously.
@@ 5.5. RQ4 - Encapsulation
Table 9. Survey questions relevant to RQ4. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parentheses on the righthand side.
ID Question Sample Distribution
9a When you use an unsafe API, how often do you look for documentation to ensure that you meet all of its requirements for safety and correctness? Used unsafe APIs 85%8%(152)
9b When you expose a safe API for unsafe code, how often do you include runtime checks to ensure that its requirements for correctness and safety are met? Exposed a safe API for unsafe code 58%29%(141)
9c When you expose an unsafe API to users, and it is their responsibility to ensure that certain requirements are met, how often do you document these requirements? Exposed an unsafe API 90%3%(68)
9d When you expose an unsafe API to other users, and it is their responsibility to ensure that certain requirements are met, how often do you include runtime checks for these requirements? 18%71%(68)
9e How often do you refactor Rust applications to remove unsafe code? All 6%88%(160)
Regularly wrote unsafe code 7%85%(85)
9f How often do you choose to avoid using an unsafe API when a safe alternative exists? All 82%10%(160)
Used unsafe APIs 83%10%(152)
Always    Most of the time    About half the time    Sometimes    Never
@@
Participants reasoned about the soundness and correctness of their unsafe code in terms of invariants. When developing operating systems and embedded applications, these invariants were provided by hardware specifications. For example, one participant who developed an embedded application would write a “magic value” (P8) to RAM to indicate which operation had triggered a reboot. The microcontroller that they used did not erase its memory on reboot, so they could rely on the value being initialized. Some participants perceived hardware-level correctness properties as “orthogonal to Rust” (P1), since they were related to the properties of the language. However, other participants who developed embedded applications relied on third-party libraries to encode hardware requirements into the type system so that their programs would fail to compile if a device had been misconfigured.
Developers leveraged Rust’s type system to uphold invariants that they believed were necessary for soundness. In operating systems, these typically began as operation-specific refinement properties on primitive types, but they quickly built up into system-level invariants; “it gets to a high-level really quickly” (P3). Rust’s aliasing restrictions were a natural fit:
@@
Another participant experimented with using ghost permissions to reason about pointer arithmetic. This method was popularized for Rust by Yanovski et al.’s GhostCell type (Yanovski et al., 2021). Other participants relied on ad hoc reasoning to ensure that their programs were correct and free of undefined behavior. Some appealed to their prior experience: “[in] C++, you do this all the time” (P15). One participant was accustomed to reasoning about aliasing rules in compiler development, so they felt confident that they were adhering to Rust’s restrictions in unsafe contexts. Other participants validated their decisions by auditing their code, but most had some level of implicit trust; “that’s sort of another… I got to trust it type thing” (P5). This aligns with Höltervennhoff et al. (Höltervennhoff et al., 2023)—a majority of their participants reasoned about unsafe code in terms of contracts or invariants, but some reported using it carelessly, and most felt that it was difficult to write correctly due to its context-sensitivity.
@@
Most survey respondents (61%) predicted that it would be at least somewhat easy for another developer at their skill level to understand a random unsafe block or function from their code. Interview participants who isolated their unsafe code had more confidence that its requirements were met, especially if they went beyond what Rust’s type system could verify. However, participants reported that unsafe code was prevalent in JIT compilers and operating systems. Domain-specific, non-local reasoning was necessary to understand any arbitrary unsafe snippet: “you have to know the innards of the JIT-compiled function…so it’s just not encapsulated” (P11). Though unsafe code may have been isolated, it is unclear if participants used it minimally. As shown in Table 0, more than 80% (9) of survey respondents only sometimes refactored their code to remove unsafe features, if at all. This distribution was similar for the subset of respondents who had regularly written unsafe code.
@@
Most interview participants and 88% of survey respondents reported that they had exposed a safe API for unsafe code. This was also common for Höltervennhoff et al. (Höltervennhoff et al., 2023), who found that 22 of their 26 participants mentioned safe interfacing. Our interview participants would create encapsulations where they perceived that the requirements for their unsafe code were satisfied by either Rust’s type system or what they reasoned to be local invariants.
However 43% of our survey respondents reported exposing unsafe APIs, relative to six of Höltervenhoff et al. (Höltervennhoff et al., 2023)’s 26 interview participants. Our participants cited the same motivations for exposing unsafe APIs as they did for using unsafe code in any capacity. One participant exposed unsafe versions of safe API endpoints which had the same behavior, but they did not include runtime checks. Another interview participant exposed unsafe APIs so that users could circumvent their safe encapsulations, in case they became difficult to use. Each of these practices aligns with another participant’s perception of a cultural value among Rust developers to make API surfaces as detailed as possible: “Rust has a tendency to… make the API as complicated as it needs to be to fully represent what is actually happening behind the scenes” (P1). Other participants indicated that they would only expose an unsafe API when it would be impossible to encapsulate without placing the burden of correctness on the user. Our survey respondents also identified with these motivations; 76% exposed unsafe APIs because it would be impossible to encapsulate them without preconditions, while 51% exposed unsafe APIs for performance and 13% did so for ergonomics. Höltervennhoff et al. (Höltervennhoff et al., 2023) also found that participants would expose unsafe APIs by necessity or to improve performance, but it was unclear if ergonomics was also a motivation.
Regardless of motivation, 61% of respondents claimed that they would only expose unsafe APIs when they had requirements beyond what Rust’s type system could verify, and they typically documented these requirements. Results for Question 9 indicate that 90% of respondents who exposed unsafe APIs would document their safety requirements most of the time, if not always. However, respondents did not typically add any additional safeguards to prevent unsafe APIs from being used incorrectly, since 71% (9) only sometimes inserted run-time checks within these APIs, if at all. This was more common when exposing safe APIs; 58% (9) inserted checks at least most of the time when exposing a safe API for unsafe code. When using unsafe APIs, 85% (9) of participants would usually look for documentation to inform their design decisions.
@@
An overwhelming majority of survey respondents also preferred safe APIs. More than 80% (9) of respondents claimed that they would choose a safe API over an unsafe alternative most of the time, if not always—regardless of whether or not they had ever used an unsafe API before.
Table 10. Survey questions relevant to RQ4. Percentages on the left include “Always” and “Most of the time”, while those on the right include “Sometimes” and “Never”. The number of respondents in each sample is included in parentheses on the righthand side.
ID Question Sample Distribution
10a When you expose a safe API for unsafe code, how often are all of its requirements for correctness and safety satisfied by the properties of Rust’s type system? Exposed a safe API for unsafe code 62%18%(141)
10b How often is the Rust community’s guidance and documentation adequate for you to know how to use unsafe correctly? All 70%16%(160)
Always    Most of the time    About half the time    Unsure    Sometimes    Never
@@
Höltervennhoff et al. (Höltervennhoff et al., 2023) had similar findings; 16 of their 26 participants felt some degree of uncertainty about whether their unsafe code was correct. Our interview participants were most often uncertain when unsafe was pervasive or when they called foreign functions. When describing their uncertainty about encapsulation, our participants compared the “theory” of safety to the “practice” of how an API would be used. When these concepts were misaligned, errors occurred. One participant observed this mismatch with a Rust crate that encapsulated a foreign library. The library used a memory-mapped file, and the API made it possible to unmap the file while the user retained a reference to it, leading to a use-after-free error. Another participant struggled with discrepancies between the interface and the implementation of Rust’s Layout API, which describes the size and alignment of an allocation. The design of Layout was changed to “tweak the safety requirements” (P14), which caused code that was previously correct to exhibit undefined behavior.
@@
A few interview participants encountered situations where the precise definition of undefined behavior was either unclear or overly restrictive. For example, one participant felt that it would be necessary to create a mutable reference to uninitialized memory. This is considered undefined behavior by Rust’s current semantics (Durham, 2022), but the participant was uncertain about how this pattern would be problematic in practice. Another participant wished to be able to take a reference to the first field of a product type and use it to create valid references to adjacent fields using integer offsets. This was considered undefined behavior under Rust’s Stacked Borrows (Jung et al., 2019) aliasing model, but it is accepted under the newer Tree Borrows model (Villani et al., 2025). Additionally, multiple participants were uncertain about how to handle the case when a program panics while a value is being dropped.
Survey respondents were somewhat more certain about their decisions. As shown in Table 0, more than 60% (10) of respondents who had exposed a safe API for unsafe code were certain at least most of the time, if not always, that its safety properties were satisfied by Rust’s type system. Additionally, 70% (10) of respondents found that the Rust community’s documentation and guidance was adequate for them to know how to use unsafe code correctly at least most of the time. However, only 6% felt that it was always adequate, suggesting that improved documentation would be helpful for the overwhelming majority of respondents. Similar to the questions in Tables 5 and  8, we added “Unsure” as a response option for the questions shown in Table 0, since they rely on developers’ subjective notions of certainty and adequacy.
@@
How do Rust developers reason about encapsulating unsafe code?